OTO AUCT ( M 6. Let f: R3 R3, f(x, y, z) = (2, x - y, y + 2). Show that f is an automorphism pol. 1 *14 Igal * Pertanyent
Answer #1

Definef: R →R’is defined by f(x,y,z)=(2, x-y,y+z) (1)For any &, BER* ,suppose f(a)= f (B) =f((X1, X1, 21)) = f(x2,42,22) (21,